### v1

@@
Guo et al.
[…]
Optimization with Multi-sourced Reference Information and Unknown Trust
[…]
Optimization with Multi-sourced Reference Information and Unknown Trust: A Distributionally Robust Approach
@@
In problems that involve input parameter information gathered from multiple data sources with varying reliability, incorporating users’ trust about different sources in decision-optimization models can potentially improve solution performance and reliability. In this work, we propose a novel multi-reference distributionally robust optimization (MR-DRO) framework, where the model inputs are uncertain and their probability distributions can be statistically inferred from multiple data sources. Via nonparametric data fusion, we construct a Wasserstein ambiguity set to minimize the worst-case expected value of a stochastic objective function, accounting for both uncertainty and unknown reliability of information sources. We reformulate the MR-DRO model as a linear program given linear objective and constraints in the original problem. We also incorporate a dynamic trust update mechanism that adjusts the trust for each source based on its performance over time. In addition, we introduce the concept of probability dominance to identify sources with dominant trust. Via solving instances of resource allocation and portfolio optimization, we demonstrate the effectiveness of the trust-informed MR-DRO approach compared to traditional optimization frameworks relying on a single data source. Our results highlight the significance of integrating (dynamic) user trust in decision making under uncertainty, particularly when given diverse and potentially conflicting input data.
@@
trust-informed decision making, distributionally robust optimization, Wasserstein ambiguity set, convex programming, probabilistic dominance
@@ 1 Introduction
Information uncertainties exist ubiquitously in real-world decision-making problems, and they often arise from limited data, imprecise measurements, and/or prediction errors of multiple data references. The existing studies and literature on stochastic or robust optimization have, respectively, assumed that one can obtain an underlying distribution or an uncertainty set of the random parameters from all data sources, before making decisions, without differentiate the quality of each data source and how they impact the quality of solutions. For example, stochastic optimization methods (see, e.g., Birge and Louveaux 2011) often assume full knowledge of the true distribution of uncertain parameters and optimize the expected objective value accordingly. In contrast, robust optimization models focus on optimizing the worst-case objective value within an uncertainty set of the random parameters (Ben-Tal and Nemirovski 2002, Bertsimas et al. 2011). Furthermore, distributionally robust optimization (DRO) (see Kuhn et al. 2024) uses statistical metrics to build an ambiguity set of the unknown probability distribution and optimize decisions considering the worst-case distribution within the set (Calafiore and Ghaoui 2006, Delage and Ye 2010, Wiesemann et al. 2014, Jiang and Guan 2016). During the past decade, the Wasserstein-metric-based ambiguity set (Kantorovich and Rubinstein 1958) has gained popularity in the DRO literature due to its convergence guarantees and computational tractability (Esfahani and Kuhn 2018, Xie 2021, Shen and Jiang 2023, Chen et al. 2024), and has been applied to a wide range of decision problems in different application contexts. Such a framework optimizes decisions by considering the worst-case performance of an objective function over an ambiguity set characterized as a Wasserstein ball of a given radius centered around an empirical distribution. However, most existing studies construct their Wasserstein ambiguity sets using empirical distributions derived from a single data source, thereby limiting their applicability in scenarios where data originate from multiple diverse sources.
In practice, data often comes from multiple information sources, making the integration of diverse reference information critical to high-quality and robust decision making. For example, weather predictions, both short- and long-term, rely on data from satellites, radar, weather stations, and ocean buoys. Similarly, medical professionals collect comprehensive health metrics such as blood pressure, heart rate, and cholesterol levels from various medical devices and health monitors to assess the health condition of a patient and make treatment decisions. A wide range of data fusion techniques have been developed to combine observations from multiple sources to form a unified system description, with applications in, for example, robotics (see Liggins II et al. 2017), healthcare (see, e.g., Albahri et al. 2023), and smart transportation (see, e.g., Ounoughi and Yahia 2023). However, to the best of our knowledge, optimization models and tools that use data from multiple sources with heterogeneous reliability have not been widely studied and explored, especially under existing uncertainties of the systems and/or environments. In a recent work, Rychener et al. (2024) propose a DRO framework to integrate heterogeneous data sources in situations where data from the true distribution are extremely scarce or unavailable. They assume that data are from multiple biased sources, each believed to be similar to but different from the true distribution. Each source distribution can then be approximated by an empirical distribution derived from the respective source data. To handle uncertainty, they build an ambiguity set by intersecting multiple Wasserstein balls, each centered at the empirical distributions from one data source. This approach mitigates the biases inherent in individual sources and improves decision making by leveraging data from multiple sources. They further demonstrate that, under specific conditions, their model can be computationally tractable and produce solutions with robust guarantees. One limitation of the results in Rychener et al. (2024) is that they assume source distributions are static and sufficiently similar to the true underlying distribution, which may not hold in cases of distribution shifts or large data discrepancies. Guo et al. (2024) tackle the challenge of managing data from multiple sources for optimization under uncertainty through a parametric data fusion method. The parametric ambiguity set is built based on samples from an estimated Normal distribution, where the first and second moments are calculated as trust-weighted sums of the corresponding moments from each source distribution. Under reasonable assumptions, the model can be reformulated as a linear program utilizing the $1$-Wasserstein metric. However, the method in Guo et al. (2024) requires all distributions involved to share the same type of distribution, e.g., Normal distribution, to ensure that the parametric data fusion process can be performed.
There exists a strong need for new methods capable of effectively integrating data from multiple sources with heterogeneous reliability into optimization models and frameworks under uncertainty. In this paper, we investigate alternative approaches that can handle diverse and potentially identify unreliable data sources. We recognize that in many practical situations, the reliability of each data source may vary and evolve over time, necessitating a mechanism to weigh and update the influence of each source dynamically. To address this challenge, we introduce the concept of “trust” to represent a decision maker’s belief in the reliability of each data source, borrowed from similar notion being considered in the literature of human-robot interaction (Lee and See 2004, Lewis et al. 2018, Wagner and Arkin 2011, Oleson et al. 2011). Incorporating trust into the data fusion process allows decision makers to weigh information according to their evaluations of each source’s historical performance, facilitating a customized approach to effectively integrate reference information from multiple sources (see more relevant discussions and justifications in e.g., Guo et al. 2024).
Moreover, the concept of trust dynamics characterizes how an individual’s trust in certain data sources evolves as more data becomes available. In the literature of human-robot interaction studies, Bayesian inference is widely used to learn parameters in personalized trust prediction models (see, e.g., Guo and Yang 2021, Guo et al. 2021). In the context of online learning, Herbster and Warmuth (1998) introduce algorithms including the static-expert and variable-share methods to dynamically allocate weights between competing experts, minimizing additional loss relative to the best expert over time. These methods provide valuable insights on how to model and leverage trust in a dynamic data fusion process.
Building on these insights, in this work, we propose a multi-reference distributionally robust optimization (MR-DRO) model that minimizes the worst-case expected cost over a Wasserstein ambiguity set constructed using a trust-weighted nonparametric data fusion approach. Our approach does not require individual sources to share the same type of distributions (Guo et al. 2024) or to be sufficiently similar to the true underlying distribution (Rychener et al. 2024). We develop a trust update framework that dynamically adjusts trust values. This mechanism evaluates the alignment of historical data between sources and true realizations, recalibrating trust to reflect the true reliability of each source. By continuously updating trust for multiple sources based on their evolving accuracy, our numerical results will show that the proposed MR-DRO approach and framework ensures that more reliable sources have greater influence on the decision-making process, enhancing the robustness and adaptability of the optimization model.
@@ 1.1 Summary of Contributions
The main contributions of this paper are threefold. First, we use a distributionally robust approach to optimize decisions under uncertainty where information about the uncertain parameters needs to be obtained from multiple data sources with different reliability. We construct the ambiguity set using nonparametric data fusion that incorporates the concept of “trust” and derive reformulation to preserve computational tractability. Second, we design a trust update process to simulate trust dynamics as additional data become available, using past errors to update trust. We introduce the concept of probability dominance to rationalize the existence of information sources with dominant trust. Third, through extensive computational experiments on stochastic resource allocation and portfolio optimization problems, we demonstrate the effectiveness of our method and validate the derived theoretical results.
@@ 1.2 Structure of the Paper
The remainder of the paper is organized as follows. In Section 2, we use the nonparametric data fusion and formulate the MR-DRO model. In Section 3, we derive tractable reformulations of the MR-DRO model. In Section 4, we introduce a trust update mechanism to simulate the evolution of trust over time and derive properties of probability dominance. In Section 5, we conduct numerical studies using instances of resource allocation and portfolio optimization, to demonstrate the computational efficiency and effectiveness of MR-DRO. In Section 6, we conclude the paper and outline potential directions for future research.
@@
Throughout the paper, we use bold symbols (i.e., $\bm{u}\in\mathbb{R}^{n}$ and $\bm{\xi}\in\mathbb{R}^{m}$) to denote the vector form of a decision variable or a parameter and use bold capital letter (i.e., $\bm{A}\in\mathbb{R}^{m\times n}$) to denote a matrix. We denote the set $\{1,\ldots,N\}$ as $[N]$. The inner product of two vectors $\bm{u},\bm{v}\in\mathbb{R}^{n}$ is expressed as $\langle\bm{u},\bm{v}\rangle:=\bm{u}^{\mathsf{T}}\bm{v}$.
@@ 2 Models and Solution Approaches
We consider a problem where a decision maker needs to integrate reference information from multiple sources with varying levels of credibility to infer uncertain parameters. We introduce the data fusion approach for handling multi-sourced reference information in Section 2.1, and present the MR-DRO model in Section 2.2.
@@ 2.1 Nonparametric data fusion for multiple sources
Let $M,H,I$ denote the dimension of uncertain parameters, the number of information sources, and the number of past events (up to the decision point), respectively. The uncertain parameters of our problem are captured by a random vector $\bm{\xi}\in\mathbb{R}^{M}$. Although the true distribution of $\bm{\xi}$ is inherently unknown, we rely on observed data points from past events to infer statistical information of $\bm{\xi}$. We use $i\in[I+1]$ to index the events as the discrete time series of data we receive from $H$ sources, where $i\in[I]$ corresponds to past events and $i=I+1$ represents the current event at which a decision needs to be made. During each past event $i\in[I]$, we receive predictions of uncertain parameters from the source $h\in[H]$ as $\tilde{\bm{\xi}}^{(i)}_{h}\in\mathbb{R}^{M}$. Following the realization of the event $i$, we observe the true values $\bm{\xi}^{(i)}_{\textrm{true}}\in\mathbb{R}^{M}$, allowing us to calculate the prediction errors of the source $h$ as $\Delta\bm{\xi}^{(i)}_{h}=\tilde{\bm{\xi}}^{(i)}_{h}-\bm{\xi}^{(i)}_{\textrm{true}}$.
@@
We assume that the predictions from each source are independent. We further assume that the prediction accuracy of each source remains consistent throughout all events, with errors from each source $h$ following an unknown distribution for all $h\in[H]$. However, it is not necessary to assume that the error distributions are identical across different sources. Therefore, utilizing the historical errors specific to each source, we can revise the predictions for the current event, and these revised predictions can serve as more reliable reference data. At the beginning of the current event $I+1$, we receive predictions $\tilde{\bm{\xi}}^{(I+1)}_{h}\in\mathbb{R}^{M}$ from source $h$. Using the errors from the past $I$ events, we compute revised predictions $\hat{\bm{\xi}}_{h}$ from source $h$ for the current event as $\hat{\bm{\xi}}_{h}=[\hat{\bm{\xi}}^{(1)}_{h},\ldots,\hat{\bm{\xi}}^{(I)}_{h}]^{\mathsf{T}}$, $\hat{\bm{\xi}}_{h}\in\mathbb{R}^{I\times M}$. Each element $\hat{\bm{\xi}}^{(i)}_{h}\in\mathbb{R}^{M}$ for all $i\in[I]$ is calculated as $\hat{\bm{\xi}}^{(i)}_{h}=\tilde{\bm{\xi}}^{(I+1)}_{h}-\Delta\bm{\xi}^{(i)}_{h}$.
@@
We design trust on different data sources as weights in the nonparametric data fusion. Specifically, we assume that a decision maker has an initial trust $\bm{t}^{(0)}=[t^{(0)}_{1},\ldots,t^{(0)}_{H}]^{\mathsf{T}},\bm{t}^{(0)}\in\mathbb{R}^{H}$ before receiving any data. The trust is updated to $\bm{t}^{(i)}$ after each event $i$ and applied in the subsequent event $i+1$, for all $i\in[I]$. We ensure that $\sum_{h=1}^{H}t^{(i)}_{h}=1$ for all $i\in[I]$ and $0\leq t^{(i)}_{h}\leq 1$ for all $i\in[I],h\in[H]$. This condition implies that the sum of the individual’s trust in each source always equals $1$, enabling the use of trust values as probabilistic weights. Consequently, we compute the distribution ^$\hat{\mathbb{P}}_{HI}$ as a discrete distribution:
^$\hat{\mathbb{P}}_{HI}(\bm{\xi}):=\sum_{h=1}^{H}\frac{t^{(I)}_{h}}{I}\sum_{i=1}^{I}\delta(\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}),$ (1)
@@
where $\hat{\bm{\xi}}^{(i)}_{h}$ corresponds to the $i$-th revised prediction from source $h$, and $\delta(\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h})$ is the Dirac delta function (Dirac 1981), which equals to $1$ when $\bm{\xi}$ matches the prediction $\hat{\bm{\xi}}^{(i)}_{h}$ and 00 otherwise.
@@
Consider a small example of a resource allocation problem in one region involving $H=2$ sources with only uncertain demand ($M=1$). Each source has provided predicted demands for $I=2$ past events such as $\tilde{\bm{\xi}}_{1}=[11,14]^{\mathsf{T}}$ and $\tilde{\bm{\xi}}_{2}=[8,14]^{\mathsf{T}}$. The realization of past events is given by $\bm{\xi}_{\text{true}}=[10,13]^{\mathsf{T}}$. With these realizations, we calculate the prediction errors for each source as $\Delta\bm{\xi}_{1}=[1,1]^{\mathsf{T}}$ and $\Delta\bm{\xi}_{2}=[-2,1]^{\mathsf{T}}$. For the current event, we update the predicted demands for each source as $\xi^{(3)}_{1}=6$ and $\xi^{(3)}_{2}=9$, respectively. Then, we compute the revised predictions from each source for the current event’s reference as $\hat{\bm{\xi}_{1}}=[5,5]^{\mathsf{T}}$ and $\hat{\bm{\xi}_{2}}=[11,8]^{\mathsf{T}}$. Suppose that we use trust $t^{(2)}_{1}=0.6$ in source 1 and $t^{(2)}_{2}=0.4$ in source 2. The empirical distribution ^$\hat{\mathbb{P}}_{HI}$ is then composed of $H\times I=4$ data points, with each point from the source $1$ having probability $0.3$, and each point from source $2$ having probability $0.2$.

### v2

@@
Optimization with Multi-sourced Information and Unknown Reliability: A Distributionally Robust Approach
@@
In problems that involve input parameter information gathered from multiple data sources with varying reliability, incorporating decision makers’ trust on different sources in optimization models can potentially improve solution performance. In this work, we propose a novel multi-reference distributionally robust optimization (MR-DRO) framework, where the model inputs are uncertain and their probability distributions can be statistically inferred from multiple information sources. Via nonparametric data fusion, we construct a Wasserstein ambiguity set to minimize the worst-case expected cost of a stochastic objective function, accounting for both uncertainty and unknown reliability of several given information sources. We reformulate the MR-DRO model as a linear program given linear objective and constraints in the original problem. We also incorporate a dynamic trust update mechanism that adjusts the trust for each source based on its performance over time. In addition, we introduce the concept of probability dominance to identify sources with dominant trust. Via computational studies using resource allocation and portfolio optimization instances, we demonstrate the effectiveness of the MR-DRO approach compared to traditional optimization frameworks relying on a single data source. Our results highlight the significance of integrating (dynamic) decision maker’s trust in optimization under uncertainty, particularly when given diverse and potentially conflicting input data.
@@
(R) Robust Optimization; multi-sourced stochastic optimization; nonparametric distributionally robust optimization; convex programming; probabilistic dominance
@@ 1 Introduction
Information uncertainty is ubiquitous in real-world decision-making problems. Such uncertainty often arises from limited data, measurement noise, and prediction errors generated by multiple information sources. Classical stochastic and robust optimization frameworks typically assume that uncertainty can be characterized a priori either by a known probability distribution or by a predefined uncertainty set constructed from available data. These approaches generally do not differentiate among data sources in terms of reliability or account for how heterogeneous data quality affects solution robustness.
In stochastic optimization \@BBOPcitep\@BAP\@BBN(Birge and Louveaux 2011)\@BBCP, the true distribution of uncertain parameters is assumed to be known, and decisions are optimized with respect to expected performance. Robust optimization \@BBOPcitep\@BAP\@BBN(Ben-Tal and Nemirovski 2002, Bertsimas et al. 2011)\@BBCP, by contrast, optimizes against the worst-case realization within a prescribed uncertainty set. Distributionally robust optimization (DRO) \@BBOPcitep\@BAP\@BBN(Kuhn et al. 2025)\@BBCP bridges these paradigms by constructing an ambiguity set of probability distributions consistent with observed data and optimizing against the worst-case distribution within that set \@BBOPcitep\@BAP\@BBN(Calafiore and Ghaoui 2006, Delage and Ye 2010, Wiesemann et al. 2014, Kuhn et al. 2024, Jiang and Guan 2016)\@BBCP. Among various constructions, Wasserstein-metric-based ambiguity sets \@BBOPcitep\@BAP\@BBN(Kantorovich and Rubinstein 1958)\@BBCP have become particularly prominent due to their statistical convergence guarantees and computational tractability \@BBOPcitep\@BAP\@BBN(Esfahani and Kuhn 2018, Xie 2021, Shen and Jiang 2023, Chen et al. 2024)\@BBCP. In most existing work, however, the Wasserstein ball is centered at a single empirical distribution derived from one data source, implicitly assuming homogeneous data quality and limiting applicability in settings with multiple heterogeneous sources.
In practice, decision-makers often rely on diverse information sources with varying levels of accuracy and bias. Weather forecasting integrates satellite, radar, and ground-station measurements; healthcare decisions synthesize signals from different diagnostic devices; and cyber-physical systems aggregate distributed sensor data. While data fusion techniques have been extensively studied in engineering and AI \@BBOPcitep\@BAP\@BBN(Liggins II et al. 2017, Albahri et al. 2023, Ounoughi and Yahia 2023)\@BBCP, their integration into optimization-based decision models under exogenous uncertainty remains limited in the OR/MS literature.
Recent DRO-based approaches begin to incorporate multi-source information. \@BBOPcitet\@BAP\@BBNRychener et al. (2024)\@BBCP construct ambiguity sets as intersections of Wasserstein balls centered at source-specific empirical distributions to hedge against source bias. Although this framework mitigates misspecification at the source level, it assumes static source distributions and sufficient proximity to the target distribution—assumptions that may fail under distribution shifts or substantial cross-source heterogeneity. \@BBOPcitet\@BAP\@BBNGuo et al. (2024)\@BBCP propose a parametric data fusion method that aggregates trust-weighted moments to form a Normal-based ambiguity set, yielding a tractable linear reformulation under the 1-Wasserstein metric. However, the reliance on a common parametric family restricts applicability when source distributions differ structurally. More recently, \@BBOPcitet\@BAP\@BBNKeehan et al. (2025)\@BBCP develop a Wasserstein DRO framework for temporally evolving multi-source data using weighted empirical distributions and concentration-based weight calibration. While this advances dynamic integration, it relies on bounded Wasserstein drift assumptions and requires nontrivial joint calibration of weights and ambiguity radii, particularly in high-dimensional settings.
These limitations reveal a fundamental gap: there is a need for optimization frameworks that (i) flexibly integrate heterogeneous data sources without restrictive parametric assumptions, (ii) explicitly model varying and evolving source reliability, and (iii) maintain tractability and robust performance guarantees.
To address this gap, we introduce a multi-reference distributionally robust optimization (MR-DRO) framework that incorporates trust as an endogenous measure of source reliability. The concept of trust, widely studied in human–robot interaction \@BBOPcitep\@BAP\@BBN(Lee and See 2004, Lewis et al. 2018, Wagner and Arkin 2011, Oleson et al. 2011)\@BBCP, captures a decision maker’s belief in the reliability of each information source. In our setting, trust governs how source data are fused into a nonparametric empirical reference distribution. Unlike existing parametric approaches, our framework does not require sources to share a common distributional family, nor does it assume proximity to a true underlying distribution.
We further develop a dynamic trust update mechanism that adjusts source weights over time based on realized performance. Drawing inspiration from Bayesian trust learning in human–robot interaction \@BBOPcitep\@BAP\@BBN(Guo and Yang 2021, Guo et al. 2021)\@BBCP and adaptive expert-weighting methods in online learning \@BBOPcitep\@BAP\@BBN(Herbster and Warmuth 1998)\@BBCP, our mechanism evaluates alignment between historical source data and observed realizations to recalibrate trust levels. This enables the model to progressively amplify reliable sources while attenuating unreliable ones.
Building on this adaptive data fusion process, the proposed MR-DRO model minimizes worst-case expected cost over a Wasserstein ambiguity set centered at a trust-weighted empirical distribution. By dynamically updating source trust and incorporating heterogeneous data in a fully nonparametric manner, the framework enhances robustness and adaptability in multi-source environments. Our numerical experiments later demonstrate that the MR-DRO approach systematically assigns greater influence to reliable sources and achieves improved out-of-sample performance compared with existing methods.
@@ 1.1 Summary of contributions
The main contributions of this paper are threefold. First, we develop a DRO approach to optimize decisions under uncertainty where information about the uncertain parameters needs to be obtained from multiple data sources that have different levels of reliability. We construct the ambiguity set using nonparametric data fusion that incorporates the concept of “trust” and derive a convex reformulation that preserves computational tractability. Second, we use a trust update process to simulate trust dynamics as additional data become available, using past errors to update trust. We introduce the concept of probability dominance to rationalize the existence of information sources with dominant trust. Third, through extensive computational experiments and numerical testing of instances of resource allocation and portfolio optimization, we demonstrate the effectiveness of our method and validate the derived theoretical results.
@@ 1.2 Structure of the paper
The remainder of the paper is organized as follows. Section 2 presents the nonparametric data fusion framework and the corresponding MR-DRO model. Section 3 develops tractable reformulations of the proposed model. Section 4 introduces the dynamic trust update mechanism, characterizes the evolution of trust across data sources over time, and establishes properties related to probability dominance. Section 5 reports numerical experiments based on resource allocation and portfolio optimization problems to evaluate the computational efficiency and performance of the MR-DRO approach. Finally, Section 6 concludes the paper and discusses directions for future research.
@@
Throughout the paper, we use bold symbols (i.e., $\bm{u}\in\mathbb{R}^{n}$ and $\bm{\xi}\in\mathbb{R}^{m}$) to denote the vector form of a decision variable or a parameter and use bold capital letters (i.e., $\bm{A}\in\mathbb{R}^{m\times n}$) to denote a matrix. We denote the set $\{1,\ldots,N\}$ as $[N]$. The inner product of two vectors $\bm{u},\bm{v}\in\mathbb{R}^{n}$ is expressed as $\langle\bm{u},\bm{v}\rangle:=\bm{u}^{\mathsf{T}}\bm{v}$.
@@ 2 Models and Solution Approaches
We introduce the data fusion approach for handling multi-sourced reference information in Section 2.1, and present the MR-DRO model in Section 2.2.
@@ 2.1 Nonparametric data fusion for multiple sources
Let $M,H,I$ denote the dimension of uncertain parameters, the number of data sources, and the number of past events (up to the decision point), respectively. The uncertain parameters of our problem are captured by a random vector $\bm{\xi}\in\mathbb{R}^{M}$. Although the true distribution of $\bm{\xi}$ is inherently unknown, we rely on observations from past events to infer statistical information of $\bm{\xi}$. We use $i$ to index the events as the discrete time series of data we receive from $H$ sources, where $i\in[I]$ corresponds to past events and $i=I+1$ represents the current event at which a decision needs to be made. At each past event $i\in[I]$, we have a prediction of the uncertainty based on each information source $h\in[H]$, denoted as $\tilde{\bm{\xi}}^{(i)}_{h}\in\mathbb{R}^{M}$. Following the realization of event $i$, we observe the true value $\bm{\xi}^{(i)}_{\textrm{true}}\in\mathbb{R}^{M}$, allowing us to calculate the prediction error of source $h$, denoted as $\Delta\bm{\xi}^{(i)}_{h}=\tilde{\bm{\xi}}^{(i)}_{h}-\bm{\xi}^{(i)}_{\textrm{true}}$, for each source $h\in[H]$.
@@
We assume that the predictions from different sources are independent. We further assume that the prediction accuracy of each source remains consistent throughout all events, with errors from each source $h$ following an unknown distribution for all $h\in[H]$. However, it is not necessary to assume that the error distributions are identical across different sources. Therefore, utilizing the historical errors specific to each source, we can revise the predictions for the current event, and these revised predictions can serve as more reliable information references. At the beginning of the current event $I+1$, we receive predictions $\tilde{\bm{\xi}}^{(I+1)}_{h}\in\mathbb{R}^{M}$ from source $h$. Using the errors from the past $I$ events, we compute revised predictions $\hat{\bm{\xi}}_{h}$ from source $h$ for the current event as $\hat{\bm{\xi}}_{h}=[\hat{\bm{\xi}}^{(1)}_{h},\ldots,\hat{\bm{\xi}}^{(I)}_{h}]^{\mathsf{T}}$, $\hat{\bm{\xi}}_{h}\in\mathbb{R}^{I\times M}$. Each element $\hat{\bm{\xi}}^{(i)}_{h}\in\mathbb{R}^{M}$ for all $i\in[I]$ is calculated as $\hat{\bm{\xi}}^{(i)}_{h}=\tilde{\bm{\xi}}^{(I+1)}_{h}-\Delta\bm{\xi}^{(i)}_{h}$.
@@
In the nonparametric data fusion, we consider trust on each data source as a weight and update the trust after each event $i\in[I+1]$. Specifically, we assume an initial trust $\bm{t}^{(0)}=[t^{(0)}_{1},\ldots,t^{(0)}_{H}]^{\mathsf{T}},\bm{t}^{(0)}\in\mathbb{R}^{H}$ before observing any data. The trust is updated to $\bm{t}^{(i)}$ after each event $i$ and applied in the subsequent event $i+1$, for all $i\in[I]$. We ensure that $\sum_{h=1}^{H}t^{(i)}_{h}=1$ for all $i\in[I]$ and $0\leq t^{(i)}_{h}\leq 1$ for all $i\in[I],h\in[H]$. Consequently, we compute the distribution $\hat{\mathbb{P}}_{HI}$ as a discrete distribution:
$\hat{\mathbb{P}}_{HI}:=\sum_{h=1}^{H}\frac{t^{(I)}_{h}}{I}\sum_{i=1}^{I}\delta_{\hat{\bm{\xi}}^{(i)}_{h}}$ (1)
@@
where $\hat{\bm{\xi}}^{(i)}_{h}$ corresponds to the $i$-th revised prediction from source $h$, and $\delta(\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h})$ is the Dirac delta function \@BBOPcitep\@BAP\@BBN(Dirac 1981)\@BBCP, which equals to $1$ when $\bm{\xi}$ matches the prediction $\hat{\bm{\xi}}^{(i)}_{h}$ and 0 otherwise.
@@
Consider a small example of a resource allocation problem in one region involving $H=2$ sources with only uncertain demand ($M=1$). Each source has provided predicted demands for $I=2$ past events such as $\tilde{\bm{\xi}}_{1}=[11,14]^{\mathsf{T}}$ and $\tilde{\bm{\xi}}_{2}=[8,14]^{\mathsf{T}}$. The realization of past events is given by $\bm{\xi}_{\text{true}}=[10,13]^{\mathsf{T}}$. With these realizations, we calculate the prediction errors for each source as $\Delta\bm{\xi}_{1}=[1,1]^{\mathsf{T}}$ and $\Delta\bm{\xi}_{2}=[-2,1]^{\mathsf{T}}$. For the current event, we update the predicted demands for each source as $\tilde{\xi}^{(3)}_{1}=6$ and $\tilde{\xi}^{(3)}_{2}=9$, respectively. Then, we compute the revised predictions from each source for the current event’s reference as $\hat{\bm{\xi}_{1}}=[5,5]^{\mathsf{T}}$ and $\hat{\bm{\xi}_{2}}=[11,8]^{\mathsf{T}}$. Suppose that we use trust $t^{(2)}_{1}=0.6$ in source 1 and $t^{(2)}_{2}=0.4$ in source 2. The empirical distribution $\hat{\mathbb{P}}_{HI}$ is then composed of $H\times I=4$ data points, with each point from the source $1$ having probability $0.3$, and each point from source $2$ having probability $0.2$.